OpenStudy (lolipop_drum312):

A hose fills a tank at a rate of 20 gallons per minute. What is the rate in liters per hour? (1 gallon ≈ 3.79 liters) Enter your answer in the box. L/h

OpenStudy (wckeller):

\[20\frac{ gallons }{ minute } * \frac{ 3.79 liters}{ 1 gallon }*\frac{ 1hour }{ 60\min }=20*3.79*\frac{ 1 }{ 60 }=1.263\frac{ liters }{ hour }\]

OpenStudy (lolipop_drum312):

Thank you

OpenStudy (wckeller):

No problem